25. Determine the length of side CB.

26.Conical containers are often used for selling popcorn or snow cones. To construct a cone, a sector is cut from a circle, then the cut edges are joined. Find the area of the sector that would be needed to make this cone.



24.0cm 18.0cm 113.1cm
А 400 700 D 15 C B
Answer #1

-> 25. Let ICBI and IABI denote the length of co and AB respectively => |ABI = 1 CB1 tant1099 and IABI=(1571CB1) tan(409) MIC=> The length of CB is 6.597 cm. 26. S 24 oli Then s=518² +24² cm = √324 +576 cm = 5900 cm = 30 cm s=30cm. => The raclius of